2015-09-10 Restart after Summer Shutdown

Mountain Manager: J Little (LBT)
OSA: G Bechetti (LBT)
Isp: E Solheid (LBT)
Observers (or, rather, Calibrators): D Thompson, O Kuhn, B Rothberg (Tucson)

*Short Summary:*
Clouds and some rain. Dave did LUCI2 calibrations, until a serious MOS error stopped things. I ran through many MODS1 calibrations. I ran into only a few hitches, among them, a couple of comm drops and one instance where "red fitsflush" was needed.

Please check the IDL license manager - I could not start IDL tonight (see IT 5671).


LUCI2:

Pupil Alignment, Field Alignment, followed by check and fine-tuning of Pupil Alignment.

Collected flexure compensation data for imaging mode at 3 elevations.

But for spectroscopic modes, no data were taken. MOS error occured when moving SpectroSieve from storage. Position of mask unknown and telescope/rotator motion stopped until further investigation by Peter or James.

MODS1:

On 6 September UT, I did some basic checks: lamps turning on, slit images, full field image, and on 7 September UT, we obtained data to measure the IMCS TTF Zeropoints. The updated tables are now in place on the mods1 computer.

Tonight will take some calibrations, with telescope at zenith and lights off in the chamber, trying to go through commonly used modes: grpixflats(dual,r,b), imflats(dual), grlamps(dual,r,b), slitflats(dual,r,b), bias3K and bias8K.

08:05: modsWake

08:05: bias8K, bias3K, bias1K

08:33: grpixflats
  • Vflat = 5: b16-20 36-46K clear, 5s, r16-20 31-38K ND1.5, 1s (brightens through the sequence)
  • Vflat = 10: b21-25 33K UG5, 2s

08:56: grpixflats_r
  • Vflat =5: r21-25 32-36.7K, ND1.5, 1s

09:14: grpixflats_b
  • Vflat = 4: b26-30, 48-55K at red end, clear, 3-sec
  • Vflat = 5: b31-35, 52K at red end, UG5, 3s

~09:35: Stop as lights go on for Elliott and Geno to access LUCI2 to turn on the MOS Web Cam

10:00: grlamps (dome lights still on)
  • grlamps_r:
    • Ne r 26 brightest line saturated, rest are not
    • XeKr r 27 ~30K peak
    • Ar r 28 ~30K peak
  • grlamps
    • Ne+Hg b 36 (27K), r 29 (60K)
    • Xe+Kr b 37 (1 line near saturation), r 30 (30K) (dome lights off now)
    • Ar b 38 (15K), r 31 (33K)
  • grlamps_b
    • Hg b 39 (30K)
    • Xe+Kr b 40 (sat) ... Above based on ds9's "view"->"horizontal graph", but the count levels reported depend on zoom factor.
    • Ar b 41 (20K)

10:25: imflats.cal
  • u b 42-46 vflat=10, 50s (20K-23K)
  • g b 47-51 vflat=4, 12s (30-34K)
  • r r 32-36 vflat=3,2s (27K)
  • i r 37-41 vflat=1.5,7s (30K)
  • z r 42-46 (24K)

11:01: slitflat_5as.cal
  • b 52-54 qth1 + ND1.5 (28K)
  • b 55-57 Qth1 + UG5 (19K)
  • r 47-49 vflat = 5 (27-29K)

11:20: slitflats.cal
  • 0.6asec: b 58-60 (30K), b 61-63 (ug5; 25K) r 50-52 (26K), r 53-55 (27K)
  • 0.8 asec: b 64-66 (28K), b 67-69 (27K), r 56-58 (22K), r 59-61 (24K)
  • 1.0 asec: b 70-72 (28K), 73-75 (30K), r 62-64 (18K), 65-67 (23K)
  • 1.2 asec: b 76-78 (30K), b 79-81 (30K), r 68-70 (21K), r 71-73 (29K)

12:42: prlamps.cal
  • prlamps_r.cal
    • Ne r 74 (1 line saturated)
    • Ar r 75 (37K max)
    • Kr r 76 (35K max)
    • Xe r 77 (50K max)
  • prlamps.cal
    • Ne+Hg r 82 (1 line saturated) b 78 (50K max)
    • Hg+Ar r 79 (37K), b (50K max)
    • Kr r 80 (42K) b 84 (40K)
    • Xe r 81 (57K) b 85 (45K)
  • prlamps_b.cal
    • Hg b 86 (59K max)
    • Ar b 87 (25K)
    • Kr b 88 (BLASTED at the red end Script used t=5-sec, clear filter, Kr only. Blue-Only does not cut off at 5650 angstroms - the lower dispersion and some bright Kr lines at lam>5650 cause this saturation. But the lower wavelength lines are OK)
    • Xe b 89 (BLASTED at the red end Script used t=20 sec, clear filter, Xe only. Same as above.)

13:10 prbias.cal

13:15 modsSleep

-- %USERSIG{OlgaKuhn - 2015-09-10}%

Comments

 
Topic revision: r3 - 10 Sep 2015, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback